OpenStudy (ksaimouli):

I-x+4I<=1 solve for x

OpenStudy (anonymous):

\[-x+4\le 1\]and\[x-4\le 1\]

OpenStudy (anonymous):

when (-x+4)>=0 ,x<=4 |-x+4|=-x+4<=1 x>=3 so 3<=x<=4(result 1) when (-x+4)<0, x>4 |-x+4|=-(-x+4)=x-4<=1 x<=5 so 4<x<=5(result 2) according to result1 and result 3<=x<5

OpenStudy (anonymous):

@JohnHanShanghai is right !

OpenStudy (anonymous):

I-x+4I<=1 -x+4 <= 1 (or) -x+4 >= -1 -x <= -3 (or) -x > = -5 x >= 3 (or) x <= 5 3<= x <= 5 note that 5 is also included in the solution

OpenStudy (anonymous):

i hate \(-x\) so i would start with the fact that \(|-x+4|=|x-4| \) i.e. start with \[|x-4|\leq 1\] \[-1\leq x-4\leq 1\] \[3\leq x\leq 5\] just a matter of taste is all

OpenStudy (anonymous):

keeps you from having to flip the inequality

OpenStudy (anonymous):

i hate -x too :)

OpenStudy (anonymous):

|-x+4|≤1 √((-x+4)^2 ) ≤1 〖(-x+4)〗^2 ≤1^2 Expanding LHS: x^2-8x+16≤1 x^2-8x+15≤0 Factorising the LHS gives (x-3)(x-5)≤0 The solution is therefore 3≤x≤5.that is values from 3 right up to 5.